OpenStudy (anonymous):

(x+5)1/2 -(5-2x)1/4=0

OpenStudy (anonymous):

Is that (x+5) "times" 1/2 or (x+5) "to the" 1/2?

OpenStudy (anonymous):

(x+5) to the 1/2

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

(x+5)^1/2 = (5-2x)^1/4 (x+5)^2 = 5-2x x^2 + 10x +25 = 5-2x x^2 +12x + 20 = 0 (x+10)(x+2)=0 so: x = -10 and x= -2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

hold on... one solution doesn't check... (see attached)

OpenStudy (anonymous):

thank you sir

OpenStudy (anonymous):

oser can you help me on another